OpenStudy (anonymous):

need help writing these expressions in radical form. please and thank you :)

OpenStudy (anonymous):

\[1) 2\frac{ 5 }{ 3 }\] 2) \[2) 4 \frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

Square it and stick it under a square-root sign.

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

\[4\times\frac{4}{3} = +\sqrt{16\times\frac{16}{9}}\]

OpenStudy (anonymous):

That's how I interpret the instructions anyway. Do you have an example to go by?

OpenStudy (anonymous):

The + sign is important because your original terms in the expressions are positive, while a square-root can be + or -.

OpenStudy (anonymous):

Ex: \[\sqrt{4} = \space ^+_- 2\] because \[ (-2)^2 = 4\] and \[(+2)^2 = 4\]